OpenStudy (anonymous):

Addison has been told that her mean average on six homework assignments in her history class is 85. She can find only five of the six assignments, which have scores of 91,72,96,88, and 75. Find the score on the lost homework. Please help me show how to solve this using the mean "x" for the missing score

OpenStudy (amistre64):

how does one go about finding a mean to start with?

OpenStudy (anonymous):

sum of all x devided by sample

OpenStudy (anonymous):

I was cooking, sorry for late reply amistre64

OpenStudy (anonymous):

@amistre64

OpenStudy (amistre64):

then lets add up the scores, divide by 6, and equate it to 85 it pretty much solves itself after that

OpenStudy (amistre64):

\[\frac{S+x}{6}=85\] \[S+x=85(6)\] \[x=85(6)-S\]

OpenStudy (anonymous):

i see, thanks a lot!

OpenStudy (amistre64):

yw

OpenStudy (anonymous):

510-422=88